OpenStudy (anonymous):

Examine the rise in gasoline prices from 2006 to 2009. The average price of regular unleaded gasoline in January 2006 was $2.31 and in January 2009 the average price of regular unleaded gasoline was $3.42. a. Use the slope formula to find the rate of change of the gasoline prices between 2006 and 2009. Explain

OpenStudy (anonymous):

The way I figure it is this, but I'm not sure if I'm right: M=y2 – y1 / x2-x1 3.42 - 2.31 / 2009 – 2006 = 1.11 / 3 = .37